Consider the standing wave pattern with 4 antinode
Answer #1

Part A

Velocity of waves on a string V=\sqrt{\frac{T}{\mu}}

There are Four antinodes.

Length of string is L=2\,m

Linear mass density \mu=3\times10^{-3}\,kg/m

since there are four antinodes, length of string is L=4\times\frac{\lambda}{2}

Hence wave length is \lambda=\frac{L}{2}=1\,m

Frequency is f=120\,Hz, Hence wave speed is V=\lambda{f}=1\times120=120\,m/s

since V=\sqrt{\frac{T}{\mu}} , T=V^{2}\mu=(120)^{2}\times3\times10^{-3}=43.2\,N

Part B

If there are n antinodes on the string, L=n\frac{\lambda}{2}=n\frac{V}{2f}

f=n\frac{V}{2L}

Using the formula V=\sqrt{\frac{T}{\mu}}

f=\frac{n}{2L}=\frac{n}{2L}\sqrt{\frac{T}{\mu}}

Frequency, linear mass density and Length are not changed. That is n\sqrt{T} does not change.

n\sqrt{T}=n'\sqrt{T'}

4\sqrt{43.2}=n'\sqrt{T'}

since Tension has increased T'>T That is n'<n.

Since the next lower integral number is 3, n'=3.

New number of antinodes is 3.

4\sqrt{43.2}=3\sqrt{T'}

T'=43.2\times\frac{4^2}{3^2}=76.8\,N